OpenStudy (anonymous):

2,3,5,8,12 A. Arithmetic B. Geormetric C. both D. neither

OpenStudy (mathmath333):

common difference here is 1,2,3,4 which is not constant so its not arithmatic sequence

OpenStudy (mathmath333):

common ratio here is 1.5,1.66,1.6,1.5 its not constant so its not geometric sequence

OpenStudy (mathmath333):

common difference =3-2=1 5-3=2 8-5=3... common ratio =3/2=1.5 =5/5=1.66...